# Nuisance-free Automatic Ground Collision Avoidance System Design: Merging Exponential-CBF and Adaptive Sliding Manifolds

### v1

@@ Abstract
The significance of the automatic ground collision avoidance system (Auto-GCAS) has been proven by considering the fatal crashes that have occurred over decades. Even though extensive efforts have been put forth to address the ground collision avoidance in the literature, the notion of being nuisance-free has not been sufficiently addressed. At this point, in this study, the Auto-GCAS design is formulated by merging exponential control barrier functions with sliding manifolds to manipulate the barrier function dynamics. The adaptive properties of the sliding manifolds are tailored to the key and governing flight parameters, ensuring that the nuisance-free requirement is satisfied. Furthermore, to ensure all safety requirements are met, a flight envelope protection algorithm is designed using control barrier functions to assess the commands generated by the Auto-GCAS. Eventually, the performance of the proposed methodology is demonstrated, focusing on authority-sharing, collision avoidance capability, and nuisance-free operation through various scenarios and Monte Carlo simulations.
Index Terms:
Safety-critical control, ground collision avoidance, flight control, authority-sharing, control barrier function
I Introduction
@@
Although the development of Auto-GCAS technology has provided significant advantages in mitigating CFIT risks, it also presents challenges stemming from human-autonomy interaction [10]. In this context, the unnecessary, untimely, and interruptive interventions of the Auto-GCAS are terminologically referred to as “nuisance” [10]. These interventions may include premature activation or false alarms during maneuvers, disrupting pilot operations and trust in the system. Moreover, interventions that are impulse-like or insufficiently aggressive may lead the pilot to perceive them as unnecessary, as they could have executed the required recovery maneuver using a more aggressive and abrupt approach [11, 12]. Therefore, to ensure harmony between human and autonomy in piloted aircraft, the design of a nuisance-free Auto-GCAS is paramount, that is why, a number of studies have been conducted in order to address this issue. Nevertheless, primarily, the qualitative and quantitative definitions of “nuisance-free” must be established. Fortunately, there is consensus on the qualitative definition: a nuisance-free intervention is characterized as “timely and aggressive” [10, 12, 13, 14, 11]. This implies that the Auto-GCAS algorithm must generate a recovery maneuver that satisfies the following criteria: (1) the applied command must have maximum allowable magnitude, such as maximum pitch rate or maximum load factor, depending on the flight control architecture, and (2) the aircraft’s closest point to the ground must match the pre-defined keep-out zone, i.e. buffer. In the quantitative sense, Eq. (1) describes the notion of nuisance-free, as clearly presented in [12].
$\begin{split}\min_{t\in[t_{0},t_{f}]}||h_{\text{A/C}}(t,\bm{u}(t))-h_{\text{DTED}}(t)||_{2}=h_{\text{buff}}\\ u_{i}(t)=\begin{cases}u_{\text{max}_{i}},\forall t\in[t_{0},t_{\text{CPA}})\\ \text{or}\\ u_{\text{min}_{i}},\forall t\in[t_{0},t_{\text{CPA}})\\ \end{cases}\end{split}$ (1)
@@
where $\bm{u}(t)$ is the control input vector. Additionally, $h_{\text{A/C}}$, $h_{\text{DTED}}$, and $h_{\text{buff}}\in\mathbb{R}$ are the altitude of aircraft, height of digital terrain elevation data (DTED), and height of pre-defined scalar buffer, respectively. Understandably, these two constraints imply that, for a given control input $\bm{u}(t)$, a timely aircraft trajectory corresponds to a recovery maneuver $h_{\text{A/C}}(t,\bm{u}(t))$ that approaches but does not violate the buffer within the time interval $t\in[t_{0},t_{f}]$, also it refers to a “last second” maneuver [12]. Furthermore, at least one control input $u_{i}(t)$ must reach its absolute maximum allowable magnitude during the interval $t\in[t_{0},t_{\text{CPA}})$, where $t_{\text{CPA}}$ is the time of the closest point of approach within $[t_{0},t_{f}]$. The illustration of nuisance and nuisance-free activations is depicted in Fig. 1.
@@
Figure 1: The illustration of maneuvers and their corresponding control inputs, which can be classified as nuisance or nuisance-free, involves $q_{\text{GCAS}}$, the pitch rate command of the Auto-GCAS, where $q_{\text{max}}$ represents its maximum magnitude.
@@
Although the primary focus is on achieving a nuisance-free design, the requirements of a complete Auto-GCAS framework extend further, necessitating adherence to the critical “do no harm” principle, which encompasses a broad range of additional considerations [10]. Fundamentally, the Auto-GCAS must command a recovery maneuver which (1) does not damage pilot or the structural integrity of the aircraft, (2) does not stimulate an uncontrollable flight, and (3) allows the pilot to intervene [10, 13, 15]. In summary, the design of a comprehensive and complete Auto-GCAS framework must integrate both the nuisance-free and “do no harm” principles, along with their associated sub-requirements. Thus, the term “complete Auto-GCAS framework” in this study refers to a design that considers these criteria.
@@
Based on the discussed challenges, there is a limited amount of unclassified research addressing the Auto-GCAS design issue. Even though the mid-air or traffic collision avoidance problem has been extensively studied—for example, for quadrotors [16], fixed-wing manned [17] and unmanned aircraft [18, 19], and even for geofencing [20]—their proposed methods cannot be directly applied to Auto-GCAS design due to the specific design criteria discussed earlier, which were not considered in those studies, with the exception of [20]. Molnar et al. [20] proposed a run-time assurance control framework for fixed-wing aircraft to ensure safety-critical tasks, such as mid-air collision avoidance and geofencing. They introduced various control barrier function (CBF) architectures, including higher-order, backstepping-based, and model-free methods, to address these issues. By merging multiple safety constraints into a unified CBF, they demonstrated the efficacy of their proposals through various simulations. However, as they highlighted as a direction for future work, the most significant limitation of their study may be the lack of additional safety constraints related to the aircraft’s safe flight envelope—e.g. measures to prevent the aircraft from exceeding the stall angle of attack—indicating that their methodology does not incorporate the “do no harm” requirements. Therefore, the available collision avoidance studies offer neither directly adoptable approaches nor a complete framework for the design of Auto-GCAS. Furthermore, in the existing accessible Auto-GCAS studies, the main focus area is generally nuisance-free design. In [13], the conceptual architecture of a nuisance-free Auto-GCAS was elaborated. The discussed architecture includes several key components, such as trajectory prediction algorithm to eliminate the potential nuisance situations. This component principally answers the question “What would the resulting trajectory look like if a recovery maneuver is initiated now?” and terrain scan pattern that scans a virtual horizon using digital terrain elevation data (DTED). Seemingly, as one of the earliest publications on nuisance-free Auto-GCAS design, the study in [13] laid a foundational framework for subsequent research, such as [21, 22, 11, 15, 14, 23, 24, 25]. As a follow-up to [13], the study in [15] also emphasized the necessity of a trajectory prediction algorithm. Similarly, the study in [11] presents a detailed framework for Auto-GCAS design, again including a trajectory prediction algorithm. Notably, several studies, including [21, 22, 23, 24], focus specifically on trajectory prediction algorithms. For instance, [21] proposed a ground collision avoidance warning and decision system with multi-trajectory risk assessment and decision functions to provide comprehensive avoidance decisions for flight crews. Likewise, [23] and [24] introduced trajectory prediction algorithms tailored for general aviation (GA). Finally, the study in [12] proposed a nuisance-free Auto-GCAS design approach by formulating timely and aggressive recovery maneuvers as an optimal control problem, thereby eliminating the need for a trajectory prediction algorithm. However, this formulation was based on simplified point-mass, three degrees-of-freedom (3DoF) flight dynamics.
As a consequence, existing Auto-GCAS literature has primarily focused on satisfying the nuisance-free criterion using computationally intensive trajectory prediction algorithms. This reliance on such algorithms often necessitates simplified kinematic 3DoF flight dynamics. Furthermore, to the best of the author’s knowledge, no unclassified study provides a complete Auto-GCAS framework that satisfies all the necessary design requirements discussed earlier.
I-A Objectives & Methodology
@@
Building on the discussed design challenges and existing literature, a pertinent research question arises: “Is there an alternative approach to guarantee nuisance-free Auto-GCAS design that reduces the computational cost of commonly used trajectory prediction algorithms while also completing the safety framework to meet all critical Auto-GCAS requirements?” To answer this, the primary objective of this study is to develop an Auto-GCAS that guarantees nuisance-free operation while adhering to stringent safety requirements, including compliance with flight envelope protection principles to meet the “do no harm” mandate. To achieve this goal, this study proposes a novel design methodology that incorporates the following key components:
@@
Additionally, a flight envelope protection algorithm using control barrier functions is designed to validate the commands generated by the Auto-GCAS, ensuring compliance with flight envelope protection constraints, including angle of attack, load factor, and bank angle. The integration of CBF/ECBF is motivated by their advantages: (1) the provision of rigorous mathematical safety guarantees, (2) the facilitation of control authority-sharing between humans and automation through their formulations, and (3) the computational efficiency through linear constrained convex optimization, enabling real-time feasibility.
@@ I-B Contributions & Organization
  • To achieve a complete safety framework, this study also designs a flight envelope protection (FEP) algorithm using control barrier functions (CBFs), which is similar to the previous work of the authors [26]. The FEP algorithm supervises Auto-GCAS commands to ensure compliance with critical flight envelope limits, such as the stall angle of attack and maximum load factor. This comprehensive framework addresses “do no harm” requirement, which is a gap observed in [20].
  • The proposed method achieves a success rate of $499$ out of $500$ Monte Carlo simulations for random initial dive cases, equating to $99.8\%$. This success rate is attributed to the study’s formally safety-proven approach, highlighting the method’s potential to significantly reduce controlled flight into terrain (CFIT) incidents.
The rest of the paper is organized as follows: In Section II, the necessary background for the study is provided, including nonlinear flight dynamics modeling and flight control law design. Section III introduces the design of the ground collision avoidance system, addressing nuisance-free design in Section III-A. To complete the safety framework, the flight envelope protection design is presented in Section IV, with specific discussions on angle of attack and load factor in Section IV-A, and bank angle in Section IV-B. Finally, the proposed architecture is rigorously assessed in Section V, through ground collision scenarios in Section V-A, pilot authority-sharing scenarios in Section V-B, and Monte Carlo simulations, including 500 different cases, in Section V-C.
II Preliminaries
@@
II-A Notations
Throughout the study, the time derivative of a $C^{1}$ function $f:\mathbb{R}^{n}\xrightarrow[]{}\mathbb{R}$ is denoted by $\dot{f}$. Moreover, a vector is denoted in the bold type, i.e. $\bm{v}$, and the vector product of two vectors $\bm{x}$ and $\bm{y}$ is denoted by $\bm{x}\times\bm{y}$, and $\text{sgn}(*)$ is the signum function. Throughout the study, the prefix $\Delta$ denotes the incremental form, i.e. $\Delta(*)$ is the incremental form of ()(*)( ∗ ). The notation of $s_{*}$, $c_{*}$, and $t_{*}$ corresponds to sine, cosine, and tangent of ()(*)( ∗ ). Finally, a control affine system is described as given in Eq. (2).
$\bm{\dot{x}}=\bm{f}(\bm{x})+\bm{g}(\bm{x})\bm{u}$ (2)
where $\bm{x}\in\mathbb{R}^{n}$ is the state vector, and $\bm{u}\in\mathbb{R}^{m}$ is the control input vector. Nonlinear mappings of $\bm{f}:\mathbb{R}^{n}\xrightarrow[]{}\mathbb{R}^{n}$ and $\bm{g}:\mathbb{R}^{n}\xrightarrow[]{}\mathbb{R}^{n\times m}$ are locally Lipschitz continuous functions.
II-B Control Barrier Functions
The superlevel set $\mathcal{C}$ is composed of the expressions in Eq. (3) by definition to be a safe set [27].
$\begin{split}\mathcal{C}=\{\bm{x}\in D\subset\mathbb{R}^{n}:b(\bm{x})\geq 0\}\\ \partial\mathcal{C}=\{\bm{x}\in D\subset\mathbb{R}^{n}:b(\bm{x})=0\}\\ Int(\mathcal{C})=\{\bm{x}\in D\subset\mathbb{R}^{n}:b(\bm{x})>0\}\end{split}$ (3)
Definition II.1.
[27] A function $b:D\subset\mathbb{R}^{n}\xrightarrow[]{}\mathbb{R}$ is the control barrier function (CBF) for the nonlinear system in Eq. (2) if the following conditions hold:
  • A superlevel set $\mathcal{C}$ exists for the function $b(\bm{x})$.
  • $b(\bm{x})$ satisfies the inequality in Eq. (4).
    $\sup_{\bm{u}\in U}\{\mathcal{L}_{f}b(\bm{x})+\mathcal{L}_{g}b(\bm{x})\bm{u}+\alpha(b(\bm{x}))\geq 0\}$ (4)
For cases where the relative degree of $\delta>1$, an extension known as the exponential control barrier function (ECBF) is employed to ensure the forward invariance property of $\mathcal{C}$.
Definition II.2.
[27, 28, 29] A smooth function $b(\bm{x})$, characterized by a relative degree $\delta$, qualifies as an exponential control barrier function (ECBF) if $\exists\kappa\in\mathbb{R}^{\delta}$ such that $\forall\bm{x}\in\mathcal{C}$,

### v2

# Nuisance-free Ground Collision Avoidance Design with Adaptive Exponential Control Barrier Functions

@@ Abstract
The significance of the automatic ground collision avoidance system (Auto-GCAS) has been proven by considering the fatal crashes that have occurred over decades. Even though extensive efforts have been put forth to address the ground collision avoidance in the literature, the notion of being nuisance-free has not been sufficiently addressed. In this study, the Auto-GCAS design is formulated by merging exponential control barrier functions with sliding manifolds to manipulate the barrier function dynamics. The adaptive properties of the sliding manifolds are tailored to the key and governing flight parameters, ensuring that the nuisance-free requirement is satisfied. Furthermore, to ensure all safety requirements are met, a flight envelope protection algorithm is designed using control barrier functions to assess the commands generated by the Auto-GCAS. Eventually, the performance of the proposed methodology is demonstrated, focusing on authority-sharing, collision avoidance capability, and nuisance-free operation through various scenarios and Monte Carlo simulations. Simulation results demonstrate that the proposed adaptive exponential CBF-based Auto-GCAS achieves a $99.88\%$ ground collision avoidance success rate across diverse scenarios, without nuisance activations and while respecting aircraft dynamic limits.
1 Introduction
@@
Although the development of Auto-GCAS technology has provided significant advantages in mitigating CFIT risks, it also presents challenges stemming from human-autonomy interaction [10]. In this context, the unnecessary, untimely, and interruptive interventions of the Auto-GCAS are terminologically referred to as “nuisance" [10]. These interventions may include premature activation or false alarms during maneuvers, disrupting pilot operations and trust in the system. Moreover, interventions that are impulse-like or insufficiently aggressive may lead the pilot to perceive them as unnecessary, as they could have executed the required recovery maneuver using a more aggressive and abrupt approach [11, 12]. Therefore, to ensure harmony between human and autonomy in piloted aircraft, the design of a nuisance-free Auto-GCAS is paramount, that is why, a number of studies have been conducted in order to address this issue. Nevertheless, primarily, the qualitative and quantitative definitions of “nuisance-free" must be established. Fortunately, there is consensus on the qualitative definition: a nuisance-free intervention is characterized as “timely and aggressive" [10, 12, 13, 14, 11]. This implies that the Auto-GCAS algorithm must generate a recovery maneuver that satisfies the following criteria: (1) the applied command must have maximum allowable magnitude, such as maximum pitch rate or maximum load factor, depending on the flight control architecture, and (2) the aircraft’s closest point to the ground must match the pre-defined keep-out zone, i.e. buffer. In the quantitative sense, Eq. (1) describes the notion of nuisance-free, as clearly presented in [12].
[…]
@@
where $\bm{u}(t):\mathbb{R}\to\mathbb{R}^{\text{m}}$ is the control input vector. Additionally, $h_{\text{A/C}}(t,\bm{u}(t)):\mathbb{R}\times\mathbb{R}^{\text{m}}\to\mathbb{R}$, $h_{\text{DTED}}(t):\mathbb{R}\to\mathbb{R}$, and $h_{\text{buff}}\in\mathbb{R}_{>0}$ are the altitude of aircraft, height of digital terrain elevation data (DTED), and height of pre-defined scalar buffer, respectively. Understandably, these two constraints imply that, for a given control input $\bm{u}(t)$, a timely aircraft trajectory corresponds to a recovery maneuver $h_{\text{A/C}}(t,\bm{u}(t))$ that approaches but does not violate the buffer within the time interval $t\in[t_{0}\hskip 4.26773ptt_{f}]$, also it refers to a “last second" maneuver [12]. Furthermore, at least one control input $u_{i}(t)$ must reach its absolute maximum allowable magnitude during the interval $t\in[t_{0}\hskip 4.26773ptt_{\text{CPA}})$, where $t_{\text{CPA}}$ is the time of the closest point of approach within $t_{0}\hskip 4.26773ptt_{f}]$. The illustration of nuisance and nuisance-free activations is depicted in Fig. 1.
@@
Figure 1: The maneuvers and control inputs, classified as nuisance or nuisance-free, involve $q_{\text{GCAS}}$, the Auto-GCAS pitch rate command limited by $q_{\text{max}}$.
@@
Although the primary focus is on achieving a nuisance-free design, the requirements of a complete Auto-GCAS framework extend further, necessitating adherence to the critical “do no harm" principle, which encompasses a broad range of additional considerations [10]. Fundamentally, the Auto-GCAS must command a recovery maneuver which (1) does not damage pilot or the structural integrity of the aircraft, (2) does not stimulate an uncontrollable flight, and (3) allows the pilot to intervene [10, 13, 15]. In summary, the design of a comprehensive and complete Auto-GCAS framework must integrate both the nuisance-free and “do no harm" principles, along with their associated sub-requirements. Thus, the term “complete Auto-GCAS framework" in this study refers to a design that considers these criteria.
1.1 Related Work
@@
Based on the discussed challenges, there is a limited amount of unclassified research addressing the Auto-GCAS design issue. Even though the mid-air or traffic collision avoidance problem has been extensively studied—for example, for quadrotors [16], fixed-wing manned [17] and unmanned aircraft [18, 19], and even for geofencing [20]—their proposed methods cannot be directly applied to Auto-GCAS design due to the specific design criteria discussed earlier, which were not considered in those studies, with the exception of [20]. Molnar et al. [20] proposed a run-time assurance control framework for fixed-wing aircraft to ensure safety-critical tasks, such as mid-air collision avoidance and geofencing. They introduced various control barrier function (CBF) architectures, including higher-order, backstepping-based, and model-free methods, to address these issues. By merging multiple safety constraints into a unified CBF, they demonstrated the efficacy of their proposals through various simulations. However, as they highlighted as a direction for future work, the most significant limitation of their study may be the lack of additional safety constraints related to the aircraft’s safe flight envelope—e.g. measures to prevent the aircraft from exceeding the stall angle of attack—indicating that their methodology does not incorporate the “do no harm" requirements. Therefore, the available collision avoidance studies offer neither directly adoptable approaches nor a complete framework for the design of Auto-GCAS. Furthermore, in the existing accessible Auto-GCAS studies, the main focus area is generally nuisance-free design. In [13], the conceptual architecture of a nuisance-free Auto-GCAS was elaborated. The discussed architecture includes several key components, such as trajectory prediction algorithm to eliminate the potential nuisance situations. This component principally answers the question “What would the resulting trajectory look like if a recovery maneuver is initiated now?" and terrain scan pattern that scans a virtual horizon using digital terrain elevation data (DTED). Seemingly, as one of the earliest publications on nuisance-free Auto-GCAS design, the study in [13] laid a foundational framework for subsequent research, such as [21, 22, 11, 15, 14, 23, 24, 25]. As a follow-up to [13], the study in [15] also emphasized the necessity of a trajectory prediction algorithm. Similarly, the study in [11] presents a detailed framework for Auto-GCAS design, again including a trajectory prediction algorithm. Notably, several studies, including [21, 22, 23, 24], focus specifically on trajectory prediction algorithms. For instance, [21] proposed a ground collision avoidance warning and decision system with multi-trajectory risk assessment and decision functions to provide comprehensive avoidance decisions for flight crews. Likewise, [23] and [24] introduced trajectory prediction algorithms tailored for general aviation (GA). Finally, the study in [12] proposed a nuisance-free Auto-GCAS design approach by formulating timely and aggressive recovery maneuvers as an optimal control problem, thereby eliminating the need for a trajectory prediction algorithm. However, this formulation was based on simplified point-mass, three degrees-of-freedom (3DoF) flight dynamics.
Consequently, most existing Auto-GCAS approaches in the literature rely on computationally intensive trajectory prediction algorithms to satisfy the nuisance-free criterion. This reliance typically limits implementations to simplified 3DoF kinematic models, restricting their applicability to realistic flight dynamics. Moreover, these methods often lack the flexibility to adapt to varying flight conditions or enforce dynamic constraints in real time. To the best of the authors’ knowledge, there is currently no unclassified study that offers a complete Auto-GCAS framework capable of simultaneously addressing nuisance avoidance, dynamic feasibility, and real-time adaptability within a unified and certifiable control structure.
1.2 Objectives & Methodology
@@
Building on the discussed design challenges and existing literature, a pertinent research question arises: “Can a computationally efficient alternative to trajectory prediction-based Auto-GCAS be developed that not only guarantees nuisance-free behavior but also provides a complete safety framework meeting all critical operational and dynamic constraints?" To answer this, the primary objective of this study is to develop an Auto-GCAS that guarantees nuisance-free operation while adhering to stringent safety requirements, including compliance with flight envelope protection principles to meet the “do no harm” mandate. To achieve this goal, this study proposes a novel design methodology that incorporates the following key components:
@@
Additionally, a flight envelope protection algorithm using control barrier functions is designed to validate the commands generated by the Auto-GCAS, ensuring compliance with flight envelope protection constraints, including angle of attack, load factor, and bank angle. The integration of CBF/ECBF is motivated by their advantages: (1) the provision of rigorous mathematical safety guarantees, (2) the facilitation of control authority-sharing between humans and automation through their formulations, and (3) the computational efficiency through linear constrained convex optimization, enabling real-time feasibility. Therefore, in contrast to trajectory planning or prediction-based approaches, CBFs require only the current state and allow online enforcement of safety without full and/or partial horizon planning. The use of exponential CBFs further enhances responsiveness near constraint boundaries, and their combination with adaptive manifolds enables tuning of intervention aggressiveness based on pitch angle, bank angle, and true airspeed.
@@ 1.3 Contributions & Organization
  • To achieve a complete safety framework, this study also designs a flight envelope protection (FEP) algorithm using control barrier functions (CBFs), which is similar to the previous work of the authors [26]. The FEP algorithm supervises Auto-GCAS commands to ensure compliance with critical flight envelope limits, such as the stall angle of attack and maximum load factor. This comprehensive framework addresses “do no harm" requirement, which is a gap observed in [20].
  • The proposed method achieves a success rate of $849$ out of $850$ Monte Carlo simulations for random initial dive cases, equating to $99.88\%$. This success rate is attributed to the study’s formally safety-proven approach, highlighting the method’s potential to significantly reduce controlled flight into terrain (CFIT) incidents.
The rest of the paper is organized as follows: In Section 2, the necessary background for the study is provided, including nonlinear flight dynamics modeling and flight control law design. Section 3 introduces the design of the ground collision avoidance system, addressing nuisance-free design in Section 3.1. To complete the safety framework, the flight envelope protection design is presented in Section 4, with specific discussions on angle of attack and load factor in Section 4.1, and bank angle in Section 4.2. Subsequently, the proposed architecture is rigorously assessed in Section 5, through ground collision scenarios in Section 5.1, pilot authority-sharing scenarios in Section 5.2, and Monte Carlo simulations, including $850$ different cases, in Section 5.3. Finally, a brief discussion is provided in Section 6 to elaborate on the potential of the proposed method.
2 Preliminaries
@@
2.1 Notations
A vector is denoted in the bold type, i.e. $\bm{v}$, and the vector product of two vectors $\bm{x}$ and $\bm{y}$ is denoted by $\bm{x}\times\bm{y}$, and $\text{sgn}(*)$ is the signum function. Throughout the study, the prefix $\Delta$ denotes the incremental form, i.e. $\Delta(*)$ is the incremental form of ()(*). The notation of $s_{*}$, $c_{*}$, and $t_{*}$ corresponds to sine, cosine, and tangent of ()(*). Finally, a control affine system is described as given in Eq. (2).
$\bm{\dot{x}}=\bm{f}(\bm{x})+\bm{g}(\bm{x})\bm{u},$ (2)
where $\bm{x}\in\mathbb{R}^{n}$ is the state vector, and $\bm{u}\in\mathbb{R}^{m}$ is the control input vector. Nonlinear mappings of $\bm{f}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ and $\bm{g}:\mathbb{R}^{n}\to\mathbb{R}^{n\times m}$ are locally Lipschitz continuous functions.
2.2 Control Barrier Functions
To formally define safety constraints in control systems, let define the notion of a barrier function $b(\bm{x})$. The function $b(\bm{x}):D\subset\mathbb{R}^{n}\to\mathbb{R}$ is designed such that a safe set $C$ can be defined as:
$C=\{\bm{x}\in D\subset\mathbb{R}^{n}\mid b(\bm{x})\geq 0\},$ (3)
where $b(\bm{x})\geq 0$ ensures that the system state remains within the safe set, $\partial C=\{x\in D\mid b(\bm{x})=0\}$ represents the boundary, and $\text{Int}(C)=\{x\in D\mid b(\bm{x})>0\}$ denotes the interior of the safe set.
Given a continuously differentiable function $b(\bm{x})$ and a dynamical system in Eq. (2), the Lie derivatives of $b(\bm{x})$ along the vector fields $\bm{f}(\bm{x})$ and $\bm{g}(\bm{x})$ are defined as:
$\mathcal{L}_{f}b(\bm{x})=\frac{\partial b(\bm{x})}{\partial\bm{x}}\bm{f}(\bm{x}),\quad\mathcal{L}_{g}b(\bm{x})=\frac{\partial b(\bm{x})}{\partial\bm{x}}\bm{g}(\bm{x}).$ (4)
The relative degree of $b(\bm{x})$ with respect to the control input $\bm{u}$ is the number of times it must be differentiated along the system dynamics before $\bm{u}$ explicitly appears. For a function $b(\bm{x})$ with relative degree $\delta$, its higher-order Lie derivatives are given by:
$\mathcal{L}_{f}^{\delta}b(\bm{x})=\frac{d^{\delta}}{dt^{\delta}}b(\bm{x}),$ (5)
where $\mathcal{L}_{f}^{\delta}b(\bm{x})$ denotes the Lie derivative of $b(\bm{x})$ to the power of its relative degree.
Definition 2.1 (Exponential Control Barrier Function).
A continuously differentiable function $b(\bm{x}):D\subset\mathbb{R}^{n}\to\mathbb{R}$ is an exponential control barrier function (ECBF) if there exists a coefficient gain vector $\kappa\in\mathbb{R}^{\delta}$ such that for all $\bm{x}\in C$: